Use our class áat to do the foloving problems Be nest and o se our class data to do the f Do the hypothesis tests on separate paper. Due at or before 1. (Ch.11) Using our class data: test the claim that the proportion of Cabrillo students who are dog owners is greater than proportion for testing are satisfied here. of Cabrillo students who are cat owners. Assume that all requirements 5 points) 2. (Ch.12.1) A research study from 1990 classified eye color into 5 categories. The study produced the following breakdown of eye colors: (5 points) 32% blue/grey rises 15% blue/grey/green irises with brown/yellow specks 12% green-light brown irises with minimal specks 16% brown irises with specks 25% dark brown irises For our purposes, let's condense this to the following: green/hazel 12% brown 41% blue 47% Complete the following table using our class data Grn/Hzl Blue Brown Eve color Test at the 1% significance level whether our class distribution of eye color the study. percentages into expected outcomes by multiplying each percentage by the sample size of our class. matches the distribution of Discuss any adjustments or choices you made with the data. Hint: Convert the given 5 points) 3. (Ch.12.2) Complete the following table from our class data. Eve color Blue Green/Hazel Brown Hair blond Color black brown Test the claim that eye color is dependent on hair color. Use a 2.5% significance level. Discuss any adjustments or choices you made with the data.
Height Age hair color Eye Color lave a Cat? lave a Dog? Work 10+ Smoker? Drink #Se ings GPA #Colege Units "of Hours #of Hors #Hons #Hours completed Studying Browsing Watching TV Exercising (inches) Hours? Alcohol? of Aloohol 1 65 19 brown blue y 2 63 20 brown brown n 3 71 19 black brown n 4 77 23 brown hazely 5 74 19 blond blue 6 67 20 blond blue 7.5 0 2.8 7 61 18 blond greeny 8 64 20 brown bley 9 68 21 black browny 10 66 19 brown blue 11 66 22 brown blue 12 62 25 blackbrown y 13 62 18 brown browny 14 58 21 black brown n 15 72 19 black green n 16 75 20 black brown n 17 68 18 black brown n 18 70 22 black brown y 19 70 19 brown blue 20 76 19 brownbrown n 21 70 33 brown brown n 22 74 20 brown brown y 23 77 22 brown brown n 24 64 22 brownbrown n 25 71 19 red 26 65 31 blue brown n 0 3.3 2 2.7 9.5 2.5 12 3.01 7.5 9.5 2 maybe n 0 inconsistent 1 2.5 0 3.69 4 3.6 0 3.2 1 3.3 7.5 3.2 0 3.8 12.5
Answer #1

1:

Out of 26 students, 12 own a cat. Out of 26 students, 12 own a dog.

Here we have x1-12, nı =26,P. = x1 / n =0.4615 x2 -12,n2 -26, P22/n-0.4615 Hypotheses are α =0.05 Level of significance Test